OpenStudy (anonymous):

if the measure of angle T is (123-5x)degrees, what is the measure of it complement

OpenStudy (lgbasallote):

let the complement be y (123 - 5x) + y = 90 therefore the complement is just y = 90 - (123 - 5x) y = 90 - 123 + 5x y = -33 + 5x y = 5x - 33

OpenStudy (lgbasallote):

got it @fielkool ?